OpenStudy (anonymous):

A ratio of right-handed to left-handed students is 9:2. There are 352 students in the school. How many students are right-handed students are there. How many left handed students are there?

OpenStudy (anonymous):

you'll need to solve this system: \(\huge \frac{9}{2}=\frac{R}{L} \) and R + L = 352 can you take it from here?

OpenStudy (anonymous):

no can you help me thank you so much

OpenStudy (anonymous):

\(\huge \frac{9}{2}=\frac{R}{L} \rightarrow 9L=2R \rightarrow \frac{9L}{2}=R\) so we'll substitute this expression for R into the second equation: R + L = 352 \(\large (\frac{9L}{2})+L=352 \) \(\large \color {red}2\cdot [(\frac{9L}{2})+L]=\color {red}2\cdot 352 \) \(\large 9L+2L=704 \) can you take it from here?

OpenStudy (anonymous):

yah i got it but how can there be 35.2 left handed people.....there is .2 of a person???